### v2

# Singularities of base loci on abelian varieties

Abstract.
We prove that the log canonical threshold of the base ideal of a complete linear system on a complex abelian variety is $\geq 1$, and equality holds if and only if the base locus has divisorial components. Consequently the same assertions hold for the ideal of the intersection of translates of theta divisors by the points of a finite subgroup.
The author is partially supported by the MIUR Excellence Department Project MatMod@TOV awarded to the Department of Mathematics of the University of Rome Tor Vergata, and by PRIN 2022 ”Moduli spaces and Birational Geometry”. He is a member of GNSAGA - INDAM
A well known theorem of Kollár ([K1, Theorem 17.13]) asserts that for a complex principally polarized abelian variety the pair $(A,\Theta)$ is log-canonical. This simple and fundamental result has been subsequently sharpened by Ein and Lazarsfeld in [EL], and there are still important open conjectures in this field, see e.g. [CGS], [MP, §H.29], [SY, §9] and references therein. In turn the theorems of Kollár and Ein-Lazarsfeld have been extended to other integral and/or $\mathbb{Q}$-divisors on abelian varieties in [H1],[H2],[DH],[P], [B], [JL].
In this note Kollár’s result is extended in a different direction: base loci of complete linear systems, and intersections of translates of theta divisors by the points belonging to a finite subgroup. We work over $\mathbb{C}$. For an effective line bundle $L$ on an abelian variety $A$ we denote $\mathfrak{b}_{L}$ the base ideal sheaf of the complete linear system $|L|$, namely the image of the twisted evaluation map $H^{0}(A,L)\otimes L^{-1}\rightarrow{\mathcal{O}}_{A}$. Our result is
Theorem.
$lct(\mathfrak{b}_{L})\geq 1$ and equality holds if and only if the zero locus of $\mathfrak{b}_{L}$ has divisorial components.
(When there are divisorial components the situation is completely understood, see the last part of the proof below). By a standard argument it follows that the conclusion of Theorem holds also for the (scheme-theoretic) intersection of translates of theta divisors of a principally polarized abelian variety $A$ by all points of a finite subgroup $H\subset A$.
Here and in the sequel, given a morphism $f:X\rightarrow Y$ and an ideal sheaf $\mathfrak{a}$ on $Y$ we will denote $f^{-1}\mathfrak{a}$ the inverse image ideal sheaf on $X$, i.e. the ideal sheaf generated by the pullbacks of local sections of $\mathfrak{a}$.
Corollary.
Let $A$ be a principally polarized abelian variety, $\Theta$ a theta divisor, and $H$ a finite subgroup of $A$. Let
$\mathfrak{a}_{\Theta,H}=\sum_{h\in H}t_{h}^{-1}{\mathcal{O}}_{A}(-\Theta)\subset{\mathcal{O}}_{A}.$
Then
$lct(\mathfrak{a}_{\Theta,H})\geq 1$
and equality holds if and only if the zero locus of $\mathfrak{a}_{\Theta,H}$ has divisorial components.
When there are no divisorial components we don’t expect our results to be close to sharp, and a better understanding of the log canonical threshold (and more generally of the jumping numbers) of base ideals on abelian varieties seems to be an interesting and difficult problem.
The motivation for this question comes from the attempt to show that $\mathfrak{b}_{L}$ is a radical ideal sheaf, i.e. the base scheme of $L$ is reduced. Or at least, as conjectured by Debarre in [D, Conjecture 2], that the 2-codimensional components of the base scheme are generically reduced (see the erratum in [ACP]). Although our result does not imply this, we hope that the method will be helpful in future developments.
The result of this paper was obtained while I was working with Enrico Arbarello and Giulio Codogni on the reducedness of the base scheme of line bundles on abelian varieties, and on the other topics of the erratum in [ACP]. I thank them for their generosity, and for valuable conversations and suggestions. More recently, I have greatly benefitted of a continuous exchange with Arbarello.
1. Background material on the Fourier-Mukai transform and generic vanishing
1.1. Notation
Given an abelian variety $A$ we denote $g=\dim A$. Moreover we denote $\widehat{A}={\rm Pic}^{0}A$ and $\mathcal{P}$ the Poincaré line bundle on $A\times\widehat{A}$. Given a point $\alpha\in\widehat{A}$ we denote $P_{\alpha}$ the corresponding line bundle on $A$. All sheaves appearing in the sequel are coherent sheaves.
Given a sheaf $\mathcal{F}$ on $A$ its cohomological support loci are
$V^{i}(A,\mathcal{F})=\{\alpha\in{\rm Pic}^{0}A\>|\>h^{i}(A,\mathcal{F}\otimes P_{\alpha})>0\}.$
1.2. (Symmetric) Fourier-Mukai transform associated to the Poincarè line bundle
Denoting $D(X)$ the bounded derived category of the category of coherent sheaves on a projective variety $X$, the Fourier-Mukai functor $D(A^)\Phi^{A\rightarrow\widehat{A}}_{\mathcal{P}}$$\rightarrow D(\widehat{A})$ (introduced in seminal paper [M]) is an exact equivalence. Its quasi-inverse is $\Phi^{\widehat{A}\rightarrow A}_{\mathcal{P}^{\vee}[g]}$.
For our purposes a version of this equivalence, introduced by Schnell in [S], will be more practical. The symmetric Fourier-Mukai transform is defined as the contravariant functor
$FM_{A}:D(A)\rightarrow D(\widehat{A})^{op},\quad FM_{A}=\Phi^{A\rightarrow\widehat{A}}_{\mathcal{P}}\circ\Delta_{A}$
where $\Delta_{A}:D(A)\rightarrow D(A)^{op}$ is the dualizing functor $\Delta_{A}={\mathbf{R}}\mathcal{H}om(\>\cdot\>,{\mathcal{O}}_{A}[g])$. Also $FM_{A}$ is an exact equivalence, with quasi inverse $FM_{\widehat{A}}$ ([S, Theorem 1.1]).
Let $\pi:A\rightarrow B$ be a surjective homomorphism with connected fibers of abelian varieties and let $i:\widehat{B}\rightarrow\widehat{A}$ be the dual inclusion. By [S, Proposition 1.1] the following properties hold:
$FM_{A}\circ\pi^{*}=i_{*}\circ FM_{B}$ (1.1)
$FM_{B}\circ Li^{*}=R\pi_{*}\circ FM_{A}.$ (1.2)
1.3. Vanishing conditions
Following [S], we have the following vanishing conditions on sheaves on abelian varieties.
(a) A sheaf $\mathcal{F}$ on $A$ is said to be a GV sheaf (i.e. satisfies generic vanishing) if $FM_{A}(\mathcal{F})$ is a sheaf (in degree zero). If this is the case we denote
$FM_{A}(\mathcal{F})=\widehat{\mathcal{F}}$ (1.3)
It follows if $\mathcal{F}$ is a GV sheaf on $A$ then also $\widehat{\mathcal{F}}$ is a GV-sheaf on $\widehat{A}$ and $\widehat{\widehat{\mathcal{F}}}$$\mathcal{F}$.
If $\mathcal{F}$ is a GV sheaf then the fibre of $\widehat{F}$ at a point $\alpha\in\widehat{A}$ is $H^{0}(A,\mathcal{F}\otimes P_{-\alpha})^{\vee}$ ([S, (10)]. It follows that the (reduced structure of the) support of $\widehat{\mathcal{F}}$ is the subvariety $V^{0}(A,\mathcal{F})$.
(b) $\mathcal{F}$ is said to be M-regular if $\mathcal{F}$ is GV and $\widehat{\mathcal{F}}$ is a torsion free sheaf. Since the support of $\widehat{\mathcal{F}}$ is $V^{0}(A,\mathcal{F})$, if $\mathcal{F}$ is M-regular then $V^{0}(A,\mathcal{F})=\widehat{A}$, i.e. $h^{0}(A,F\otimes P_{\alpha})>0$ for all $\alpha\in\widehat{A}$.
(c) $\mathcal{F}$ is said to be an IT0 sheaf (i.e. satisfies the index theorem with index 0) if $\mathcal{F}$ is GV and $\widehat{\mathcal{F}}$ is a locally free sheaf.
Although not directly used in the sequel, it is perhaps useful to keep in mind the following well known characterizations of the above conditions in terms of the cohomological support loci (which are more usually taken as definitions in the literature):
A coherent sheaf $\mathcal{F}$ is GV (respectively M-regular, IT0) if and only if $\operatorname{codim}_{\widehat{A}}V^{i}(A,\mathcal{F})\geq i$ (resp. >i>i, empty) for all $i>0$.
For the GV condition this is [PP1, Corollary 3.12] (but one direction was proved earlier in [H]). For the M-regularity condition this is [PP2, Proposition 2.8] and for the IT(0) condition in one direction it follows easily from base change and in the other one it follows e.g. from [PP1, Remark 3.13].
1.4. Chen-Jiang decompositions and theorem
Chen-Jiang decompositions were introduced by Z. Jiang and J. Chen in the paper [CJ, Theorem 1.1]. A sheaf $\mathcal{F}$ on an abelian variety $A$ is said to admit a Chen-Jiang decomposition if
$\mathcal{F}\cong\bigoplus_{i}(\pi_{B_{i}}^{*}\mathcal{F}_{i})\otimes P_{\alpha_{i}}$
where: $\pi_{b_{i}}:A\rightarrow B_{i}$ are surjective homomorphisms with connected fibres, $\mathcal{F}$ are M-regular sheaves on $B_{i}$, and $\alpha_{i}\in\widehat{A}$ are points of finite order.
Some remarks about this notion:
(a) Chen-Jiang decomposition are essentially unique ([LPS, Lemma 3.5] and comments after Proposition 3.3, or [CLP, §9]).
(b) Unless $\dim B_{i}=\dim A$ (i.e. the homomorphism with connected fibres $\pi_{i}$ is an isomorphism), the pullbacks of M-regular sheaves $\pi_{B_{i}}^{*}\mathcal{F}_{i}$ are GV but not M-regular (in fact by (1.1) $FM_{A}(\pi_{B_{i}}^{*}\mathcal{F}_{i})=i_{\widehat{B_{i}}*}\widehat{F_{i}}$ is a torsion sheaf).
(c) Therefore a sheaf $\mathcal{F}$ admitting a Chen-Jiang decomposition is always GV, but it is M-regular if and only if the decomposition has $\mathcal{F}$ itself as the only summand.
(d) A direct summand of a sheaf admitting a Chen-Jiang decomposition admits a Chen-Jiang decomposition ([S, Proposition 3.6]).
The theorem of Chen-Jiang [CJ, Theorem 1.1] (together with the subsequent generalization in [PPS]) summarizes and strengthens a good deal of the generic vanishing theorems of Green-Lazarsfeld and Hacon ([GL1],[GL2]), [H]). It states that:
If $f:X\rightarrow A$ is a generically finite morphism from a smooth variety to an abelian variety then the sheaf $f_{*}(\omega_{X})$ admits a Chen-Jiang decomposition.
2. Proofs
2.1. A preliminary lemma.
Let $L$ be an effective line bundle on an abelian variety $A$. Let $K(L)$ be the subgroup $\{x\in A\>|\>t_{x}^{*}L\cong L\}$. It acts on $A$ by translations. The line bundle $L$ is ample if and only if $K(L)$ is a finite group.
Lemma 2.1.1.
Let $L$ be an ample line bundle on an abelian variety $A$. Let $\mathcal{I}\subset{\mathcal{O}}_{A}$ be a $K(L)$-invariant ideal sheaf, i.e. $t_{x}^{-1}\mathcal{I}=\mathcal{I}$ for all $x\in K(L)$. Then, for $y\in A$, either $H^{0}(A,t_{y}^{*}L\otimes\mathcal{I})=H^{0}(A,t_{y}^{*}L)$ or $H^{0}(A,t_{y}^{*}L\otimes\mathcal{I})=0$. Moreover if $\mathcal{I}$ is non-trivial then the Zariski open set of $y\in A$ such that $H^{0}(A,t_{y}^{*}L\otimes\mathcal{I})=0$ is non-empty.
Proof.
We briefly recall, somewhat informally (we refer to the Mumford’s paper for the appropriate invariant treatment), the following fundamental construction, introduced in [Mu]. Let $\mathcal{G}(L)$ be the theta-group of $L$, namely the group of pairs $(x,\varphi)$, where $x\in K(L)$ and $\varphi:L\rightarrow t_{x}^{*}L$ is an isomorphism. This theta-group sits in a central extension
$0\rightarrow\mathbb{C}^{*}\rightarrow\mathcal{G}(L)\rightarrow K(L)\rightarrow 0.$ (2.1)
The group $K(L)$ does not act naturally on the global sections space $H^{0}(A,L)$ but $\mathcal{G}(L)$ does, in such a way that $\mathbb{C}^{*}$ acts with weight $1$, i.e. by homoteties. Mumford’s theorem ([Mu, §1, Proposition 3]) states that, up to isomorphism, this is the only irreducible representation of $\mathcal{G}(L)$ where $\mathbb{C}^{*}$ acts with weight $1$. Of course $K(t_{y}^{*}L)=K(L)$ for all points $y\in A$, and the extensions (2.1) for $L$ and $t_{y}^{*}L$ are naturally isomorphic, as well as the representations $H^{0}(A,L)$ and $H^{0}(A,t_{y}^{*}L)$. If $\mathcal{I}$ is a $K(L)$-invariant ideal sheaf then, for all $y\in A$, $H^{0}(A,t_{y}^{*}L\otimes\mathcal{I})$ is a representation of $\mathcal{G}(L)$ where $\mathbb{C}^{*}$ acts with weight one, and $h^{0}(A,t_{y}^{*}L\otimes\mathcal{I})\leq h^{0}(A,L)$ . Therefore by Mumford’s theorem either $h^{0}(A,t_{y}^{*}L\otimes\mathcal{I})$ equals $h^{0}(A,L)$ or it is zero. If $\mathcal{I}$ is non-trivial then $h^{0}(A,t_{y}^{*}L\otimes\mathcal{I})$ must be zero for some $y$, otherwise the zero scheme of $\mathcal{I}$ would be contained in the base locus of all translates of $L$. ∎
2.2. Proof of the Theorem
Given a rational number $c>0$ we denote $\mathcal{J}(\mathfrak{b}_{L}^{c})$ the multiplier ideal sheaf associated to $c$ and $\mathfrak{b}_{L}$. We refer to [L] for generalities about multiplier ideal sheaves.
Theorem Theorem states that $\mathcal{J}(\mathfrak{b}_{L}^{c})$ is trivial as soon as $c<1$ and that $\mathcal{J}(\mathfrak{b}_{L})$ is trivial if and only if the zero locus of $\mathfrak{b}$ has no divisorial components.
In order to prove this, in the first place we can assume that the effective line bundle $L$ is ample. Indeed otherwise, as it is well known, there is a surjective homomorphism with connected fibers $\pi:A\rightarrow B$ and an ample line bundle $M$ on $B$ such that $L=\pi^{*}M$ (one takes $B=A/K_{0}(L)$, where $K_{0}(L)$ is the neutral component of $K(L)$). Therefore $H^{0}(A,L)=\pi^{*}H^{0}(B,M)$, hence $\mathfrak{b}_{L}=\pi^{-1}\mathfrak{b}_{M}$. Thus $\mathcal{J}(\mathfrak{b}_{L}^{c}))=\pi^{-1}\mathcal{J}(\mathfrak{b}_{M}^{c})$ ([L, Example 9.5.45]).
The proof of the first assertion of the Theorem is standard. Let $L$ be an ample line bundle and $c<1$. By Nadel vanishing $V^{i}(L\otimes\mathcal{J}(\mathfrak{b}_{L}^{c})=\emptyset$ for $i>0$. Therefore the sheaf $L\otimes\mathcal{J}(\mathfrak{b}_{L}^{c})$ is IT0 and the transform $\yhwidehat{L\otimes\mathcal{J}(\mathfrak{b}_{L}^{c})}$ is locally free of rank
$h^{0}(A,L\otimes\mathcal{J}(\mathfrak{b}_{L}^{c})\otimes P_{\alpha}):=h^{0}.$
Let us recall that, for all $x\in A$, $t_{x}^{*}L\cong L\otimes P_{\alpha}$ for some $\alpha\in\widehat{A}$. If $\mathcal{J}(\mathfrak{b}_{L}^{c})$ was not trivial then $h^{0}=0$, because the zero locus of $\mathcal{J}(\mathfrak{b}_{L}^{c})$ cannot not be contained in the base locus of all translates of $L$. Therefore $\yhwidehat{L\otimes\mathcal{J}(\mathfrak{b}_{L}^{c})}=0$, hence $L\otimes\mathcal{J}(\mathfrak{b}_{L}^{c})=0$. Since this is impossible $\mathcal{J}(\mathfrak{b}_{L}^{c})$ is trivial.
Let us turn to the second assertion of the Theorem. Assume that $\mathcal{J}(\mathfrak{b}_{L})$ is non-trivial. We will prove that then the zero locus of $\mathfrak{b}_{L}$ has divisorial components. To begin with, we recall that: There exists a smooth projective variety $X$, and a generically finite morphism $f:X\to A$ such that $L\otimes\mathcal{J}(\mathfrak{b}_{L})$ is a direct summmand of $f_{*}\omega_{X}$. Let us review this. In fact
$\mathcal{J}(\mathfrak{b}_{L})=\mathcal{J}(\frac{1}{k}E)$
where $k>1$ and $E=E_{1}+\cdots+E_{k}$, where the $E_{i}$’s are general divisors in $|L|$, see [L, Prop. 9.2.22, and Prop. 9.2.26]. Thus $E$ is a divisor in $|kL|$ such that $\frac{1}{k}E]=0$, and it follows from the work of Esnault-Viehweg ([EV, (3.13)]) that the sheaf $\mathcal{J}(\frac{1}{k}E)\otimes L$ is a direct summand of the pushforward of a canonical bundle. A very readable account of this is [DH, Case $t=1$ p.226].
Thus by the theorem of Chen-Jiang plus Remark (d) of Subsection 1.4 the (indecomposable!) sheaf $\mathcal{J}(\frac{1}{k}E)\otimes L$ has a Chen-Jiang decomposition. Therefore: there exist a surjective homomorphism of abelian varieties $\pi_{B}:A\rightarrow B$ with connected fibers, a M-regular sheaf $\mathcal{G}$ on $B$, and a finite order point $\beta\in\widehat{A}$ such that
$\mathcal{J}(\mathfrak{b}_{L})\otimes L\cong P_{\beta}\otimes\pi_{B}^{*}\mathcal{G}.$ (2.2)
The base ideal $\mathfrak{b}_{L}$ is $K(L)$-invariant. Hence also the multiplier ideal $\mathcal{J}(\mathfrak{b}_{L})$ is $K(L)$-invariant, because, according [K2, §3.3] one can take a $K(L)$-invariant log resolution of $\mathfrak{b}$. Therefore by Lemma 2.1.1 $H^{0}(A,\mathcal{J}(\mathfrak{b}_{L})\otimes L\otimes P_{\alpha})=0$ for $\alpha$ general in $\widehat{A}$. Hence the GV sheaf $J(\mathfrak{b}_{L})\otimes L$, is not M-regular (item (b) of Subsection 1.3). Therefore $\dim B<\dim A$ in (2.2).
Since $\mathcal{G}$ is a M-regular sheaf on $B$ we have that $V^{0}(B,\mathcal{F})=\widehat{B}$ and therefore it follows from (2.2) and projection formula that
$V^{0}(A,L\otimes\mathcal{J}(\mathfrak{b}))=i_{\widehat{B}}(\widehat{B})-\beta$
[…]
$\mathcal{J}(\mathfrak{b}_{L})\otimes L\cong\pi_{B}^{*}\mathcal{G}.$ (2.3)
Next, we claim that
$\mathcal{G}\cong{\pi_{B}}_{*}L\>.$ (2.4)
To prove this, we apply the symmetric Fourier-Mukai transform to (2.3). Using (1.1) we have
$\yhwidehat{\mathcal{J}(\mathfrak{b}_{L})\otimes L}={i_{\widehat{B}}}_{*}\widehat{\mathcal{G}}$ (2.5)
(where $\widehat{\mathcal{G}}$ is understood to be $FM_{B}(\mathcal{G})$). On the other hand, we claim that from Lemma 2.1.1 it follows that
$\yhwidehat{\mathcal{J}(\mathfrak{b}_{L})\otimes L}={i_{\widehat{B}}}_{*}i_{\widehat{B}}^{*}\widehat{L}\>$. (2.6)
Indeed applying the functor $FM_{A}$ to the inclusion $\mathcal{J}(\mathfrak{b})\otimes L\hookrightarrow L$ one gets a surjection
$\widehat{L}\rightarrow\yhwidehat{\mathcal{J}(\mathfrak{b})\otimes L}\rightarrow 0.$ (2.7)
[…]
Comparing (2.6) and (2.5) we get
$\widehat{\mathcal{G}}\cong i_{\widehat{B}}^{*}\widehat{L}$ (2.8)
Finally, applying the functor $FM_{\widehat{A}}$ to (2.8) and using (1.2) we get what claimed, i.e. (2.4).
Finally, plugging (2.4) into (2.3) we get
$L\otimes\mathcal{J}(\mathfrak{b}_{L})\cong\pi_{B}^{*}{\pi_{B}}_{*}L$
Since $\pi_{B}^{*}{\pi_{B}}_{*}L$ is locally free, $\mathcal{J}(\mathfrak{b}_{L})$ must be a line bundle. Since $\mathcal{J}(\mathfrak{b}_{L})$ is assumed to be non-trivial, $\mathfrak{b}_{L}$ has some divisorial components. This proves one direction of the second assertion of the Theorem.
Conversely, if $L$ is ample and $\mathfrak{b}_{L}$ has some divisorial components it is known that $A\cong B\times C$, where $B$ is a principally polarized abelian variety, $C$ is an abelian variety, and $L={\mathcal{O}}_{B}(\Theta_{B})\boxtimes L_{C}$, where $L_{C}$ is an ample line bundle on $C$ without base divisors ([BL, Theorem 4.3.1]). It follows that $\mathfrak{b}_{L}={\mathcal{O}}_{B}(-\Theta_{B})\boxtimes\mathfrak{b}_{L_{C}}$. Hence, by the above, $\mathcal{J}(\mathfrak{b}_{L})=p_{B}^{*}{\mathcal{O}}_{B}(-\Theta_{B})$. This concludes the proof of the Theorem.
2.3. Proof of the Corollary
We prove more generally the following
Corollary 2.3.1.
Let $L$ be an effective line bundle on an abelian variety and let $H$ be a finite subgroup of $A$. Then
$lct(\sum_{h\in H}t_{h}^{-1}\mathfrak{b}_{L})\geq 1$
and equality holds if and only if the zero scheme of $\sum_{h\in H}t_{h}^{-1}\mathfrak{b}_{L}$ has some divisorial components.
Proof.
The argument is standard and similar to [ACP, Corollary 8]. Let $\varphi_{L}:A\rightarrow{\rm Pic}^{0}A$ be the homomorphism associated to $L$ and let $K:=\varphi_{L}(H)$. Let us consider the isogeny
$f:{\rm Pic}^{0}A\rightarrow({\rm Pic}^{0}A)/K$
and let
$\pi_{K}:B\rightarrow A$
be the dual isogeny. Then
$H^{0}(B,\pi_{K}^{*}L)\cong\bigoplus_{\alpha\in K}\pi_{K}^{*}H^{0}(A,L\otimes P_{\alpha}).$
Therefore
$\mathfrak{b}_{\pi_{K}^{*}L}=\pi_{K}^{-1}(\sum_{\alpha\in K}\mathfrak{b}_{L\otimes P_{\alpha}})=\pi_{K}^{-1}(\sum_{h\in H}t_{h}^{-1}\mathfrak{b}_{L}).$
Corollary 2.3.1 follows from the Theorem because $\mathcal{J}(\mathfrak{b}_{\pi_{K}^{*}L})=\pi_{K}^{-1}\mathcal{J}(\sum_{h\in H}t_{h}^{-1}\mathfrak{b}_{L})$ ([L, Example 9.5.44]). ∎